### v2

# Time-IMM: A Dataset and Benchmark for Irregular Multimodal Multivariate Time Series

Abstract
Time series data in real-world applications such as healthcare, climate modeling, and finance are often irregular, multimodal, and messy, with varying sampling rates, asynchronous modalities, and pervasive missingness. However, existing benchmarks typically assume clean, regularly sampled, unimodal data, creating a significant gap between research and real-world deployment. We introduce Time-IMM, a dataset specifically designed to capture cause-driven irregularity in multimodal multivariate time series. Time-IMM represents nine distinct types of time series irregularity, categorized into trigger-based, constraint-based, and artifact-based mechanisms. Complementing the dataset, we introduce IMM-TSF, a benchmark library for forecasting on irregular multimodal time series, enabling asynchronous integration and realistic evaluation. IMM-TSF includes specialized fusion modules, including a timestamp-to-text fusion module and a multimodality fusion module, which support both recency-aware averaging and attention-based integration strategies. Empirical results demonstrate that explicitly modeling multimodality on irregular time series data leads to substantial gains in forecasting performance. Time-IMM and IMM-TSF provide a foundation for advancing time series analysis under real-world conditions. The dataset is publicly available at https://www.kaggle.com/datasets/blacksnail789521/time-imm/data, and the benchmark library can be accessed at https://anonymous.4open.science/r/IMMTSF_NeurIPS2025.
1 Introduction
Time series analysis plays a foundational role across diverse fields such as healthcare healthcare_1 ; healthcare_2 , climate science latent_ode ; llm4ts , and finance fnspid ; informer . It underpins critical applications such as patient health monitoring, weather forecasting, disaster response coordination, and financial trend prediction. These applications increasingly require modeling time series data that exhibit irregular sampling, asynchronous observation patterns, and pervasive missingness, often across multiple modalities. However, traditional time series research has largely relied on clean, unimodal, regularly-sampled datasets where observations arrive at fixed intervals. These assumptions rarely hold in real-world scenarios. In practice, time series data are often irregular, multimodal, and messy, with variable sampling rates, asynchronous modalities, and pervasive missingness.
Despite the importance of addressing such irregularities, most publicly available time series benchmarks largely overlook these complexities. Datasets such as the UCR archive ucr_archive , M4 m4 , and many multimodal datasets like Time-MMD time_mmd assume regular temporal grids and fixed sampling intervals. They fail to reflect the diverse irregularities arising in operational systems, human behavior, and natural processes. This lack of realism limits the development and evaluation of models capable of handling the challenges faced in real-world deployments.
The primary obstacle to advancing time series analysis under real-world conditions lies in the absence of benchmarks that accurately reflect the irregular, heterogeneous nature of practical data. Despite the growing body of research on both time series forecasting and multimodal learning, three critical challenges remain: (1) The oversight of real-world irregularity in existing time series benchmarks, as most prior datasets time_mmd ; m4 ; ucr_archive prioritize clean, regularly-sampled settings and fail to capture the pragmatic complexities of operational systems and natural processes; (2) The limited exploration of multimodal integration within irregular time series modeling, where existing work t_patchgnn ; latent_ode ; cru ; neural_flows ; mtand primarily focuses on unimodal numerical sequences and neglects the asynchronous nature of textual and other auxiliary modalities; and (3) The missing systematic understanding of irregularity causes, as prior efforts rarely distinguish between the major categories of irregularity — such as trigger-based events, constraint-based sampling restrictions, and artifact-based system failures — that underlie observed irregular patterns, limiting causal interpretability and model robustness.
To address these challenges, we introduce a new dataset and benchmark library designed for systematic study of irregular multimodal multivariate time series analysis. Our main contributions are:
  • Cause-Driven Irregular Multimodal Dataset. We construct Time-IMM, the first benchmark that not only explicitly categorizes real-world irregularities into nine cause-driven types—spanning trigger-based, constraint-based, and artifact-based mechanisms—but also incorporates richly annotated textual data alongside numerical observations. The textual modality captures asynchronous, auxiliary information such as clinical notes, sensor logs, or event descriptions, providing critical context for interpreting and forecasting multivariate time series. This enables comprehensive modeling of complex, cross-modal interactions and irregular sampling behaviors across multiple domains.
  • Benchmark Library for Irregular Multimodal Time Series Forecasting. We develop IMM-TSF, a plug-and-play benchmarking library for forecasting on irregular multimodal time series. It supports asynchronous integration of numerical and textual data through modular components for encoding and fusion, enabling flexible and realistic experimentation.
  • Empirical Validation of Irregularity and Multimodality Modeling. Through extensive experiments across diverse types and causes of irregularity, we show that incorporating multimodality on top of irregular time series modeling yields substantial forecasting improvements—achieving an average MSE reduction of 6.71%, and up to 38.38% in datasets with highly informative textual signals. These results underscore the superiority of multimodal approaches in handling the complexities of real-world time series forecasting.
By introducing Time-IMM and IMM-TSF, we aim to catalyze research in robust, generalizable time series analysis, moving the field beyond simplified assumptions toward truly real-world scenarios. Further details on related work and dataset limitations are provided in Appendix A and Appendix B, respectively.
2 Time-IMM: A Dataset for Irregular Multimodal Multivariate Time Series
While prior benchmarks in time series forecasting and multimodal learning have expanded the diversity of application domains, they largely overlook the intrinsic irregularities that characterize real-world data collection. Existing datasets predominantly assume regular sampling patterns, simple temporal alignments, and domain-driven organization, abstracting away the diverse causes that drive irregular observations multi_modal_forecaster ; mtbench ; chattime . Even recent multimodal datasets time_mmd focus on domain variety rather than addressing the deeper methodological challenge of irregularity. As a result, current benchmarks fail to capture the pragmatic complexities introduced by trigger-based events, constraint-driven sampling, and artifact-induced disruptions. To bridge this gap, Time-IMM explicitly centers irregularity as a core characteristic of time series data, categorizing its causes and curating multimodal datasets that reflect real-world irregular structures.
2.1 Taxonomy of Irregularity in Time Series
Irregularities in real-world time series are driven by diverse and systematic causes, not just random or arbitrary occurrences. Understanding the origins of these irregularities is crucial for developing models that are robust to realistic conditions and capable of meaningful interpretation. To structure this complexity, we propose a taxonomy based on the primary nature of the factors causing irregularity. An overview of our taxonomy is illustrated in Figure 2.
First, Trigger-Based Irregularities arise when data collection is initiated by external events or internal system triggers, meaning observations occur only in response to specific happenings rather than at predetermined intervals. Second, Constraint-Based Irregularities are driven by limitations on when data can be collected, such as operational schedules, resource availability, or human factors, where the sampling process is shaped by external restrictions rather than an ideal or intended schedule. Third, Artifact-Based Irregularities result from imperfections in the data collection process itself, including failures, delays, and asynchrony, where the intended regularity is disrupted by technical or systemic artifacts.
Figure 1: Overview of the taxonomy of irregularities in time series data. Each type is exemplified by a corresponding dataset curated as part of the Time-IMM dataset.
Figure 2: Visualization of three representative features from each dataset in Time-IMM, illustrating the diverse sampling patterns and distinct types of irregularity.
This taxonomy highlights that real-world irregularities are not caused by a single factor but arise from fundamentally different sources, each requiring tailored modeling strategies to capture their structure and implications. Each dataset in Time-IMM exemplifies a distinct type of irregularity, as shown in Figure 2. To complement the taxonomy, Figure 2 presents representative visualizations of three features per dataset, illustrating the diversity of sampling patterns and irregular behaviors. Table 1 summarizes dataset statistics in Time-IMM. We also introduce a set of metrics to characterize time series irregularity—feature observability entropy, temporal observability entropy, and mean inter-observation interval—which respectively quantify: (1) the variability of missingness across features via the normalized Shannon entropy of the feature-presence distribution, (2) the dispersion of events over time using the normalized Shannon entropy of the event-time distribution, and (3) the average time gap between successive observations. Formal definitions of these metrics are provided in Appendix C.
Table 1: Overview of datasets in Time-IMM. Columns marked with $\dagger$ refer to the textual modality.
Dataset # Entities # Features
# Unique
Timestamps
# Observations
Feature
Observability
Entropy
Temporal
Observability
Entropy
Mean
Inter-Obseration
Interval
# Text
[…]
Textual Temporal
Observability
[…]
GDELT 8 5 34317 193205 1 0.9964 7.2364 hours 14357 0.9896
RepoHealth 4 10 6783 67830 1 0.9658 1.8217 days 12310 0.9821
MIMIC 20 30 91098 219949 0.8461 0.6556 14.6157 minutes 1593 0.6758
FNSPID 10 6 3659 209688 1 0.9969 1.4507 days 20826 0.9488
ClusterTrace 3 11 12615 69001 0.893 0.9753 18.1425 minutes 688 0.9971
StudentLife 20 9 1743 153610 0.92 0.9775 1.0191 days 6623 0.9761
ILINet 1 11 4918 4918 0.9267 1 6.989 days 650 1
CESNET 30 10 51107 512760 1 1 1.17 hours 224 0.9869
EPA-Air 8 4 6587 49552 0.3777 0.9576 1.0242 hours 1244 0.9956
2.1.1 Trigger-Based Irregularities
Trigger-based irregularities are characterized by observations that occur only in response to discrete events or system triggers. We further categorize this group into three representative types, each illustrating different real-world mechanisms leading to event-driven sampling patterns.
Event-Based Logging captures scenarios where observations occur only when notable external events arise, such as geopolitical protests or seismic activity. To represent this pattern, we collect event-driven time series from the GDELT Global Database gdelt_paper , where each entry corresponds to a discrete global event like a political crisis or major policy change.
Adaptive or Reactive Sampling describes cases where the system dynamically adjusts its sampling rate based on observed activity levels, such as environmental monitoring systems increasing sampling during severe weather or networks logging more frequently during high congestion periods. We illustrate this in our benchmark with RepoHealth, a dataset collected from GitHub project histories where sampling intervals adaptively shorten during periods of high commit activity and lengthen during quieter phases.
Human-Initiated Observations involve data recording triggered by human judgment or manual action, such as doctors measuring patient vitals based on clinical necessity or users reporting symptoms only when they perceive significant changes. To capture this behavior, we include samples from the MIMIC clinical database mimicIV ; MimicNote , where vital signs like heart rate and blood pressure are recorded irregularly according to clinicians’ assessments of patient needs.
2.1.2 Constraint-Based Irregularities
Constraint-based irregularities occur when limitations on resources, schedules, or human factors restrict when data can be collected, shaping the sampling pattern according to external constraints rather than an intended schedule. We further categorize this group into three representative types that reflect how operational and resource factors introduce structured irregularities.
Operational Window Sampling refers to situations where observations are possible only during predefined operational periods, such as stock exchanges being open only during business hours or satellites collecting imagery only during orbital passes. To represent this type of irregularity, we construct datasets from FNSPIDfnspid , where financial time series are naturally sampled only during market operating hours, leading to regular gaps outside trading periods.
Resource-Aware Collection describes cases where data is sampled conditionally based on system activity, often to reduce overhead or avoid redundant logging. In large-scale compute clusters, for example, telemetry may be collected only while jobs are running, leading to activity-dependent irregularity. We illustrate this with the ClusterTrace datasetclustertrace , which captures GPU cluster activity at Alibaba over eight weeks. Here, machine-level metrics are recorded only when containers are active, producing structured gaps tied to workload presence and resource scheduling.
Human Scheduling / Availability involves data collection governed by human activity schedules, such as smartphone sensor logs being recorded primarily when users are awake, or clinical measurements being taken during scheduled patient rounds rather than continuously. To illustrate this phenomenon, we include data from the StudentLifestudentlife study, where smartphone sensing data is inherently irregular due to human behavioral patterns and daily schedules.
2.1.3 Artifact-Based Irregularities
Artifact-based irregularities arise when technical imperfections, failures, or system noise disrupt the intended regularity of data collection, creating unintended gaps or distortions in the observation sequence. We further categorize this group into three representative types that reflect different sources of system-induced irregularity.
Missing Data / Gaps occur when expected data points are not recorded, leaving empty time steps in the sequence. These gaps often result from incomplete submissions, system errors, or data quality issues. In public health settings, missing values can arise when clinics do not submit reports or when technical problems affect data collection. We use the ILINet dataset time_mmd from the CDC to demonstrate this type of irregularity. While ILINet provides weekly records of influenza-like illness across the U.S., some weeks contain missing or partial data due to non-reporting providers or processing issues.
Scheduling Jitter / Delay happens when data points are recorded at uneven time intervals due to system delays or timing issues. Instead of arriving on a regular schedule, the data shows random gaps and timing shifts. This can be caused by logging processes competing for resources, background maintenance tasks, or data throttling during peak hours. We use the CESNET dataset cesnet to show this kind of irregularity since its network flow records often appear at inconsistent times due to internal scheduling and system constraints.
Multi-Source Asynchrony arises when observations from multiple data streams are integrated despite differing internal clocks, sampling rates, or synchronization policies. Examples include environmental sensor networks where temperature, humidity, and air quality sensors report measurements at unsynchronized intervals, and health monitoring systems where wearable devices and clinical records log data at different temporal granularities. To represent this form of irregularity, we compile datasets from EPA Outdoor Air Sensors epa_air_quality , where heterogeneous sensors report at varying frequencies, creating inherent asynchrony across measurement streams.
2.2 Dataset Construction
To construct Time-IMM, we start with real-world time series data exhibiting diverse forms of irregularity, then follow a two-stage pipeline: curating relevant textual data to pair with each time series, and preserving asynchronous timestamps across both modalities to reflect real-world conditions.
Multimodal Textual Data Curation.
Each time series in Time-IMM is paired with contextual textual information, such as situational narratives, observational reports, or human-generated logs, depending on the nature of the dataset. Since the modality and format of text differ across sources, we adopt dataset-specific strategies for acquisition and preprocessing. Details on data sources and collection pipelines for each dataset are provided in Appendix D.
To ensure the textual modality is both relevant and informative, we apply a preprocessing strategy inspired by Time-MMD time_mmd . This includes an initial filtering step to retain only semantically relevant documents, followed by summarization to improve clarity and compress dense information into concise, interpretable content. Summarization is performed using GPT-4.1 Nano openai_gpt41_nano with prompt instructions tailored to preserve factual grounding, avoid hallucination, and allow abstention when relevance is uncertain. Prompt templates and preprocessing examples are included in Appendix E and Appendix F.
Handling Asynchronous Timestamps.
Unlike traditional multimodal datasets time_mmd ; multimodalmimic that assume or enforce temporal alignment between modalities, Time-IMM explicitly preserves asynchronous timestamping. Numerical time series and textual observations are allowed to occur on independent, irregular schedules. We do not impose alignment between the timestamps of numerical and textual data—this reflects real-world conditions where modalities are generated through fundamentally different processes (e.g., continuous sensors vs. episodic reports). This design choice enables more flexible modeling, supporting realistic multimodal fusion strategies that can account for both recency and content semantics.
2.3 Ethical Considerations and Data Access
All datasets in Time-IMM are derived from publicly available sources, with appropriate handling of sensitive domains such as healthcare. We ensure responsible data curation and do not redistribute any content that may violate licensing terms or privacy safeguards. Further details on ethical practices, licensing, and data access policies are provided in Appendix G.
Figure 3: Problem formulation for irregular multimodal multivariate time series forecasting.
Figure 4: IMM-TSF architecture. The library includes modular fusion layers that combine irregular numerical sequences with asynchronous text via timestamp-to-text fusion and multimodal forecast fusion.
3 IMM-TSF: A Benchmark Library for Irregular Multimodal Multivariate Time Series Forecasting
3.1 Problem Formulation
In standard irregular unimodal time series forecasting (TSF), we aim to predict future values of a numerical sequence using its past, irregularly sampled observations. Let $X=\{[(t_{i}^{(n)},x_{i}^{(n)})]_{i=1}^{P_{n}}\}_{n=1}^{N}$, where there are $N$ variables and the $n$-th variable contains $P_{n}$ past observations. Each time series observation $x_{i}^{(n)}\in\mathbb{R}$ occurs at time $t_{i}^{(n)}\in[T_{\text{start}},T_{\text{cut}}]$, where $T_{\text{start}}$ denotes the start time and $T_{\text{cut}}$ denotes the cut-off time separating past from future data. We define a set of future query times as $Q=\{[q_{k}^{(n)}]_{k=1}^{F_{n}}\}_{n=1}^{N}$, where each $q_{k}^{(n)}\in(T_{\text{cut}},T_{\text{end}}]$ and $F_{n}$ denotes the number of future queries for variable $n$. The forecasting objective is to learn a function $f_{\theta}:(X,Q)\mapsto Y$, where the predictions are given by $Y=\{[\hat{y}_{k}^{(n)}]_{k=1}^{F_{n}}\}_{n=1}^{N}$, with each $\hat{y}_{k}^{(n)}$ approximating the true time series value at time $q_{k}^{(n)}$.
To incorporate unstructured textual information, we extend the irregular TSF setup to a multimodal setting. In addition to $X$, we observe a sequence of time-stamped textual data $S=\{(\tau_{j},s_{j})\}_{j=1}^{L_{S}}$, where $s_{j}$ is a text input observed at time $\tau_{j}\in[T_{\text{start}},T_{\text{cut}}]$, and $L_{S}$ denotes the number of text observations. Here, $T_{\text{start}}$ is the shared start time for both the time series and the text modality, and no alignment is assumed between the time series timestamps $\{t_{i}^{(n)}\}$ and the text timestamps $\{\tau_{j}\}$. The multimodal forecasting objective is to learn a function $g_{\theta}:(X,S,Q)\mapsto Y$ that integrates the temporal dynamics of the irregular time series and contextual signals from the text to produce accurate predictions $Y=\{[\hat{y}_{k}^{(n)}]_{k=1}^{F_{n}}\}_{n=1}^{N}$ at the query times $Q\subset(T_{\text{cut}},T_{\text{end}}]$.
3.2 Multimodal TSF Library
To support research on forecasting under realistic multimodal irregularity, we introduce IMM-TSF, a modular benchmark library designed for plug-and-play experimentation. An overview of the library architecture is shown in Figure 4. IMM-TSF supports flexible composition of forecasting pipelines by exposing unified interfaces for numerical encoders, textual encoders, and fusion strategies. Each component is designed to be self-contained and easily swappable, allowing researchers to explore different architectural choices without needing to modify the rest of the pipeline.
Irregular Unimodal TSF Model
IMM-TSF supports both irregular time series models and regular forecasting architectures adapted to irregular settings. To adapt regular time series models for irregular data, we follow the canonical pre-alignment strategy introduced in t-PatchGNN t_patchgnn . This method interpolates irregular data to a fixed temporal grid, treating relative time and missingness masks as additional input variables. Since forecast queries remain part of the input in our problem formulation, we append their timestamps to the pre-aligned sequence during inference, enabling the model to condition predictions on the desired query times. Further implementation details are provided in Appendix H.
Text Encoder
Textual inputs are processed using pretrained language models, which are frozen during training to prevent overfitting. IMM-TSF supports a variety of open-source LLMs and compact encoders, including GPT-2 gpt2 , BERT bert , LLaMA-3.1-8B llama3 , and DeepSeek-7B deepseek . The resulting embeddings form the asynchronous textual input stream for downstream fusion.
Timestamp-to-Text Fusion (TTF) Module
In real-world settings, textual data typically arrives asynchronously and is not temporally aligned with the timestamps of numerical time series. The Timestamp-to-Text Fusion module addresses this by using the timestamps of past textual data to produce temporally aware representations for each forecast query. IMM-TSF provides two variants. (1) Recency-Weighted Averaging (RecAvg) performs a Gaussian-weighted aggregation of past text embeddings, where weights decrease with the squared time difference between the text and forecast timestamps. This allows the model to emphasize temporally proximate text while remaining simple and efficient. (2) Time2Vec-Augmented Cross-Attention (T2V-XAttn) enriches each past text embedding with a Time2Vec time2vec encoding of its timestamp, then uses cross-attention with a learned forecast query to selectively attend to temporally and semantically relevant past text.
Multimodality Fusion (MMF) Module
The Multimodality Fusion module combines text-derived representations with numerical features before prediction, allowing textual context to directly influence the final forecast. IMM-TSF includes two variants. (1) GRU-Gated Residual Addition (GR-Add) encodes a residual correction from the concatenated text and numerical features using a GRU, and blends it into the forecast via a learned gating mechanism that adapts the contribution of text dynamically. (2) Cross-Attention Addition (XAttn-Add) applies multi-head attention between numerical queries and text representations, and adds a scaled residual update to the forecast using a convex combination controlled by a fixed mixing weight. Further architectural details for both TTF and MMF modules are provided in Appendix I.
4 TSF Experiments on Irregular Multimodal Multivariate Time Series
4.1 Experimental Setup
Datasets and Forecasting Setup.
We evaluate all models on the nine Time-IMM datasets, each corresponding to a distinct cause of real-world time series irregularity. Each dataset contains multiple entities (e.g., patients, devices, repositories), and we define an entity-specific forecasting task, where both the input and target windows are sampled from the same entity without crossing entity boundaries. We apply a sliding window approach, dividing each window into a past (context) segment and a future (query) segment, enabling models to learn from historical observations and predict future values. Data is split chronologically into 60% training, 20% validation, and 20% test. Window sizes for both context and query segments are dataset-specific, reflecting native timestamp distributions and sampling patterns. Further configuration details are provided in Appendix J.
Baselines.
We compare a broad range of models categorized into three groups: (1) Regular Time Series Forecasting Models, including Informer informer , DLinear dlinear , PatchTST patchtst , TimesNet timesnet , and TimeMixer timemixer ; (2) Large Time Series Models, including the LLM-based TimeLLM timellm and the foundation model TTM ttm ; and (3) Irregular Time Series Forecasting Models, including CRU cru , Latent-ODE latent_ode , Neural Flow neural_flows , and t-PatchGNN t_patchgnn . We evaluate each model in both unimodal and multimodal settings, using mean squared error (MSE) as our primary performance metric.
Training and Optimization.
All models are trained using the Adam optimizer with a learning rate of 1e-3 and a batch size of 8. Training proceeds until early stopping based on validation loss. For model-specific hyperparameters such as the number of layers or hidden dimensions, we use the default settings provided by the original implementations. In multimodal settings, we use the best-performing combination of text encoder, Timestamp-to-Text Fusion (TTF) module, and Multimodality Fusion (MMF) module selected via validation performance. Full details for each baseline and multimodal configuration are provided in Appendix K.
Figure 5: Radar chart comparing forecasting performance across all baseline models. Shaded regions highlight the relative improvement from multimodal over unimodal variants.
4.2 Results and Analysis
Effectiveness of Multimodality.
Across all baselines, we observe consistent performance gains when incorporating textual signals alongside irregular time series data. Figure 5 shows that multimodal variants outperform unimodal counterparts in nearly every case. While prior work like Time-MMD demonstrated the value of text in regular time series forecasting, our results extend this insight to irregular settings. On average, the best multimodal configurations reduce MSE by 6.71%, with gains reaching up to 38.38% in datasets with informative text—confirming that textual context offers complementary cues under real-world irregularities. Detailed results are provided in Appendix L. Among irregular models, t-PatchGNN achieves the strongest results—consistent with its status as the state-of-the-art in irregular unimodal forecasting—and further validates the quality of our carefully curated irregular benchmark. We also find that models suited to small-scale data—such as the few-shot capable TimeLLM and TTM, and the lightweight DLinear—remain competitive in the multimodal setting. This may be due to the smaller size of multimodal datasets compared to unimodal ones, where models that are both simple and possess strong few-shot capabilities are easier to train effectively.
Impact of Irregularity Patterns.
Forecasting difficulty in the unimodal setting varies substantially across different subtypes of time series irregularity. Figure 6(a) plots per-dataset forecasting errors across baselines, grouped by the irregularity subtypes defined in our taxonomy. Datasets characterized by Missing Data (ILINet), Resource-Aware Collection (ClusterTrace), or Event-Based Logging (GDELT) are notably more difficult, as their observation patterns are sparse, erratic, or driven by latent semantic triggers that are inherently harder to model. In contrast, datasets such as RepoHealth (Adaptive Sampling), FNSPID (Operational Window Sampling), and EPA-Air (Multi-Source Asynchrony) tend to be easier to model due to more predictable or recoverable observation structures. For instance, in the FNSPID dataset, which involves stock price forecasting, the irregularity stems from known business-hour trading windows—an expected pattern that simplifies temporal reasoning.
When extending to the multimodal setting, we find that performance gains from incorporating text are not determined solely by the type of irregularity. Instead, improvements vary by dataset and are closely tied to the predictive value of the textual modality. For instance, in FNSPID, where numerical forecasting is already strong due to the regularity of stock market hours, textual news adds limited additional benefit. Conversely, ClusterTrace sees substantial gains from textual integration because its workload annotations describe the currently running processes, which are directly relevant for forecasting future resource usage. These findings highlight that the value of textual data in multimodal irregular forecasting depends more on the semantic alignment and predictive strength of the text than on the irregularity pattern itself.
Ablation of Fusion Modules.
To evaluate the design choices within our fusion framework, we conduct ablations over the Timestamp-to-Text Fusion (TTF) and Multimodality Fusion (MMF) modules, as summarized in Figure 6(b). For TTF, we observe no notable performance difference between the Time2Vec-augmented cross-attention (T2V-XAttn) and the simpler recency-based averaging (RecAvg), suggesting that both strategies for handling irregular timestamps in textual data are comparably effective. For MMF, GRU-gated residual addition (GR-Add) consistently outperforms cross-attention addition (XAttn-Add), suggesting that learnable gating provides a more effective means of integrating text-derived context into time series representations. By selectively controlling the flow of information, the gating mechanism in GR-Add helps suppress noisy or less relevant textual signals, leading to more stable and robust fusion.
Role of the Text Encoder.
We evaluate the impact of different LLM backbones on forecasting performance in the irregular multimodal TSF setting by varying the text encoder within IMM-TSF. Specifically, we test GPT-2, BERT, LLaMA 3.1 (8B), and DeepSeek (7B), and assess their performance across representative datasets. As shown in Figure 6(c), we observe no significant performance differences across these models, suggesting that the choice of LLM backbone does not strongly influence results under irregular conditions. This lack of sensitivity may reflect the nature of the irregular TSF task: because forecasting hinges more on temporal alignment and contextual anchoring than on deep semantic reasoning, increasing LLM capacity offers no clear benefit. Additionally, while our TTF and MMF modules enable integration of textual signals into the forecasting pipeline, the overall scale of our datasets remains relatively small compared to the large-scale data used to pretrained LLMs, which may limit the advantages of larger models in this setting.
Figure 6: Multimodal forecasting analysis: (a) gains across datasets; (b) fusion strategies; (c) frozen LLM backbones.
5 Future Work
Forecasting Without Known Query Timestamps.
Following prior work such as t-PatchGNN t_patchgnn , we assume that forecast query times are given, simplifying the task to predicting values at predefined future timestamps. A natural extension is to consider more open-ended settings where the model must first infer when important events or state changes are likely to occur, and then predict their values. This setup—akin to timestamp prediction or event forecasting—introduces additional uncertainty and demands joint modeling of temporal dynamics and event saliency.
Beyond Forecasting Tasks.
While IMM-TSF currently focuses on forecasting as the primary task, real-world irregular multimodal time series often require broader capabilities such as anomaly detection, classification, or retrieval. Adapting our framework to support these tasks would enable more comprehensive evaluation of modeling techniques and open new directions for multimodal irregular time series learning under operational constraints.
Beyond Textual Modalities.
Our current benchmark focuses on unstructured text as the secondary modality, but many real-world applications involve other asynchronous inputs such as images, audio, or tabular data. Extending irregular time series datasets to support multimodal integration beyond text remains an important direction for building truly general purpose time series models.
6 Conclusion
We introduce Time-IMM, the first benchmark explicitly designed around cause-driven irregularities in multimodal multivariate time series. Our taxonomy covers three distinct categories of irregularity—trigger-based, constraint-based, and artifact-based—captured across nine real-world datasets. To complement this, we present IMM-TSF, a plug-and-play library for forecasting on irregular multimodal time series, with modular support for textual encoders and fusion strategies. Empirical results show that integrating textual information improves forecasting accuracy across all irregularity types, underscoring the need for robust multimodal modeling under realistic conditions. Time-IMM and IMM-TSF are positioned to serve as strong foundations for advancing time series research toward more general and practically relevant solutions.
Appendix A Related Dataset Work
Recent progress in time series analysis has been supported by a growing number of publicly available benchmarks ts_bench_1 ; ts_bench_2 ; ts_bench_3 ; ts_bench_4 . However, most existing datasets make simplified assumptions that limit their realism and utility for modeling complex real-world phenomena. In particular, prior time series datasets often fall short along several key dimensions that Time-IMM is explicitly designed to address.
Lack of Irregularity. Many standard time series datasets assume regularly sampled data where all features are observed at fixed intervals informer ; uci_electricity_data ; caltrans_traffic_data ; ncei_weather_data . Even in multimodal contexts, existing benchmarks often maintain regular sampling, smoothing over the irregular behaviors common in real-world systems time_mmd ; ts_survey . This assumption neglects the complexities found in domains like healthcare, environmental sensing, and finance, where sampling can be triggered by events, human decisions, or operational constraints mugrep ; health_monitoring ; gru_ode .
Lack of Asynchronous Textual Modality. Most existing time series datasets include only numerical observations and lack any unstructured textual modality ucr_archive ; m4 . While some benchmarks incorporate alternative modalities such as images, audio, or video lrs3 ; wildfirets ; multimodal_survey_1 ; multimodal_survey_2 , this work focuses specifically on unstructured textual data as the complementary modality to time series. Yet in many domains, textual signals—such as clinical notes, system logs, or policy descriptions—are crucial for interpretability and context-aware forecasting log_forecasting ; clinic_note_predicting ; policy_forecasting . Even in datasets that include both time series and text time_mmd ; multimodalmimic , the modalities are often artificially aligned by enforcing shared timestamps, disregarding the fact that different modalities in real-world settings typically follow inherently asynchronous temporal patterns. Our work explicitly supports asynchronous timestamps across modalities, preserving their natural temporal structure for more realistic modeling.
Lack of Reasoning About Irregularity Causes. While some prior datasets contain irregularly sampled time series t_patchgnn ; physionet ; activity ; ushcn , they do not provide any structured explanation or annotation of why the irregularity occurs. Irregular sampling is typically treated as a nuisance to be smoothed or interpolated away, rather than a meaningful signal that could inform model design or interpretation ts_imputation_1 ; ts_imputation_2 ; ts_imputation_3 . This lack of causal reasoning limits model robustness and interpretability, especially in high-stakes domains such as healthcare, environmental monitoring, or fault diagnostics environment_monitoring ; coke . In contrast, our work introduces a structured taxonomy of irregularity, categorizing it into trigger-based, constraint-based, and artifact-based causes to support more interpretable and realistic modeling.
Appendix B Limitations
While our proposed dataset and benchmark provide a valuable foundation for studying irregular multimodal time series forecasting, several limitations remain.
First, the current version of our dataset focuses exclusively on English-language textual data. This choice simplifies preprocessing and alignment but limits the applicability of our benchmark in multilingual or cross-lingual settings, which are common in global-scale systems such as public health surveillance or international finance global_rainfall ; global_trends ; financial_predicting .
Second, our benchmark is currently limited to the task of time series forecasting. We chose this task in part because it does not require additional label collection—making it easier to construct realistic datasets from naturally occurring logs or signals timedrl ; llm4ts_small . Moreover, forecasting serves as a critical upstream task in time series analysis: models that accurately capture future dynamics can support a range of downstream applications including anomaly detection, classification, and decision support run ; ts2vec . Nonetheless, expanding the benchmark to cover these broader tasks would enable more comprehensive evaluation of modeling techniques.
Third, our focus on textual modality reflects the increasing utility of natural language as a rich, unstructured source of auxiliary information. This design decision is motivated by the recent success of large language models (LLMs), which have demonstrated strong capabilities in extracting and contextualizing semantic information from text. However, we do not currently support other modalities such as images, audio, or tabular records—despite their relevance in domains like remote sensing, clinical medicine, or human activity recognition ts_tabular ; ts_audio ; medfuse .
Fourth, we use pretrained LLMs as frozen text encoders and do not explore task-specific fine-tuning. This design simplifies benchmarking and reduces computational cost, but it may underutilize the full potential of the language models, especially in domains where domain-specific adaptation could improve alignment with time series signals calf ; news_to_forecast . To isolate the value of pretrained representations, we conducted additional experiments using Doc2Vec doc2vec as a train-from-scratch alternative. These results, presented in Appendix L.1, confirm that frozen LLMs consistently outperform Doc2Vec, reinforcing their utility even without fine-tuning.
Finally, while we introduce modular mechanisms for timestamp-aware fusion of text and time series, we do not explore more advanced strategies such as retrieval-augmented generation (RAG) timerag ; ts_rag , memory modules optimus ; tsmae , or reasoning ts_reasoning ; chatts , which could further enhance model interpretability and performance.
Despite these limitations, our work offers a structured and extensible platform for modeling real-world time series data under conditions of irregularity and multimodality, and we hope it encourages future research into more general, robust, and semantically grounded forecasting systems.
Appendix C Formalization of Irregularity Metrics
We introduce three metrics to quantitatively characterize the irregularity of time series data across entities and modalities. These metrics are designed to be interpretable, normalized where appropriate, and sensitive to structural patterns in missingness and temporal spacing.
C.1 Feature Observability Entropy
This metric quantifies how evenly features are observed across the dataset. Let $N$ be the number of features, and $f_{i}$ the number of non-missing observations for feature $i$. Define the proportion of observations for each feature as:
$p_{i}=\frac{f_{i}}{\sum_{j=1}^{N}f_{j}}$ (1)
The normalized entropy is computed as:
$\hat{H}_{\text{feat}}=-\frac{1}{\log N}\sum_{i=1}^{N}p_{i}\log(p_{i}+\epsilon)$ (2)
where $\epsilon$ is a small constant added for numerical stability.
Interpretation: A higher value of $\hat{H}_{\text{feat}}$ (closer to 1) indicates that observations are evenly distributed across all features. A lower value (closer to 0) suggests that only a subset of features dominate the observation space, revealing imbalance in feature observability.
C.2 Temporal Observability Entropy
This metric measures the dispersion of observations over time by dividing the full time span into $K$ equal-width bins and evaluating the entropy of observation counts within each bin. Let the total time range be divided into $K$ bins of equal width, and let $c_{k}$ be the number of observations falling into bin $k$. The normalized count in each bin is:
$p_{k}=\frac{c_{k}}{\sum_{j=1}^{K}c_{j}}$ (3)
The normalized temporal entropy is computed as:
$\hat{H}_{\text{temp}}=-\frac{1}{\log K}\sum_{k=1}^{K}p_{k}\log(p_{k}+\epsilon)$ (4)
where $\epsilon$ is a small constant added for numerical stability. We fix $K=10$ across all datasets to provide a stable and consistent measure of temporal dispersion.
Interpretation: A higher value of $\hat{H}_{\text{temp}}$ (closer to 1) indicates that observations are evenly distributed across time, suggesting regular sampling behavior. A lower value (closer to 0) reflects temporal clustering or bursty updates.
C.3 Mean Inter-Observation Interval (IOI)
This metric quantifies the average time gap between consecutive observations.
Let $\Delta t_{i}=t_{i}-t_{i-1}$ denote the interval between successive timestamps across all entities, and let $M$ be the total number of such intervals. The mean IOI is:
Mean $\text{Mean IOI}=\frac{1}{M}\sum_{i=1}^{M}\frac{\Delta t_{i}}{s}$ (5)
where $s$ is the number of seconds per time unit (e.g., 86,400 for days).
Interpretation: Smaller values indicate denser data (frequent sampling), while larger values suggest sparser time series. This metric is unnormalized and reported in interpretable units (e.g., hours or days).
Appendix D Dataset Details and Data Sources
We organize Time-IMM’s nine datasets according to our irregularity taxonomy. For each, we describe the data, point to the source, and summarize how numerical and textual streams were collected.
D.1 Event-Based Logging (Trigger-Based)
Description
Observations occur only when external events trigger data collection. This pattern reflects systems that respond to real-world happenings such as political crises, protests, or policy changes.
Source
GDELT Event Database 2.0 (2021–2024) gdelt_paper . Raw data was downloaded from: http://data.gdeltproject.org/gdeltv2/masterfilelist.txt
Entity
Each entity corresponds to a unique (country, event type) pair. We select 4 countries—USA, GBR, AUS, and CAN—and 2 event types: AGR (Agriculture) and LAB (Labor Union). This results in 8 distinct entities in total.
Collection
  • Numerical Time Series: We use GDELT’s event dataset from 2021/01/01 to 2024/12/31 and extract five key quantitative features per event: GoldsteinScale, NumMentions, NumSources, NumArticles, and AvgTone. Events are grouped by (country, type) to create entity-specific time series.
  • Textual Data: For each event, we download the associated news article using the source URL provided in GDELT. These articles are filtered for relevance and summarized into five-sentence narratives using GPT-4.1 Nano, following prompt instructions designed to ensure factuality, brevity, and contextual clarity.
D.2 Adaptive or Reactive Sampling (Trigger-Based)
Description
Sampling frequency is adjusted dynamically based on system activity levels. In this case, repository activity determines when state is logged—shorter intervals during development bursts and longer intervals during quiet periods.
Source
We use GitHub repository data from the GitHub Archive via Google BigQuery.
Entity
Each entity corresponds to a single GitHub repository. This results in four total entities in the dataset:
  1. 1.
    facebook/react
  2. 2.
    huggingface/transformers
  3. 3.
    kubernetes/kubernetes
  4. 4.
    pytorch/pytorch
Collection
  • Numerical Time Series: We compute a 10-dimensional numerical time series from GitHub metadata, including issue activity, pull request lifecycle statistics, backlog growth, and age metrics. The data is sampled on an adaptive schedule determined by commit frequency: 1 day during high activity, 7 days for baseline, and 14 days during low activity. Each timestamp reflects a state summary at the chosen interval.
  • Textual Data: For each sampled interval, we gather all issue and pull request comments made since the previous timestamp. We limit the number of comments to 5 per issue to control input length. These comments are aggregated by repository and date, then summarized using GPT-4.1 Nano with a prompt designed to extract key concerns, proposals, and unresolved questions. The generated summaries consist of up to five sentences and serve as the unstructured modality aligned with each adaptive sample.
D.3 Human-Initiated Observation (Event-Based)
Description
Data is recorded when humans initiate measurement or documentation, such as clinicians logging vitals based on clinical necessity. This results in inherently irregular sampling tied to human judgment and workflow.
Source
MIMIC-IV clinical database and MIMIC-Notes mimicIV ; MimicNote . Available at: https://physionet.org/content/mimiciv/3.1/
Entity
Each entity is a single ICU patient who satisfies the following inclusion criteria:
  • Patients included in the Metavision system.
  • Exactly one hospital admission (to avoid multi-stay confounds).
  • Length of stay exceeds 50 days, ensuring sufficiently long time-series.
  • Age at admission is $\geq 15$ years.
  • Has at least one chart-events record and appears in the MIMIC-IV Notes table.
From the remaining population we draw a simple random sample of $n=20$ patients to capture diverse clinical courses and recording patterns. The choice of criteria and the preprocessing code mostly follow the logic from the GRU-ODE-Bayes preprocessing pipeline: https://github.com/edebrouwer/gru_ode_bayes/tree/master/data_preproc/MIMIC
Collection
  • Numerical Time-Series
    For each selected patient we extract 105 variables spanning: IV infusion and boluses, laboratory chemistry and hematology, blood products and fluids, outputs such as urine and stool. Measurements are documented only when clinically indicated, yielding patient-specific, irregular sampling.
  • Textual Data
    All free-text notes (e.g. progress, nursing, radiology) linked to the same patients are retained in raw form together with their original timestamps. These notes provide asynchronous narrative context that can precede, coincide with, or follow numerical measurements.
This protocol produces a multimodal dataset—irregular physiological signals paired with temporally aligned clinical narratives—tailored for longitudinal modelling tasks.
D.4 Operational Window Sampling (Constraint-Based)
Description
This form of irregularity occurs when observations are limited to predefined operational hours, such as trading windows on financial markets. No data is collected during weekends, holidays, or outside business hours, resulting in structured temporal gaps.
Source
FNSPID: Financial News & Stock Prices fnspid .
Entity
Each entity corresponds to an individual publicly traded company. We first filter out companies that lack recorded time stamps or related articles, and then randomly sample 10 different companies.
Collection
  • Numerical Time Series: We download daily closing prices and related trading statistics for each selected company from https://huggingface.co/datasets/Zihan1004/FNSPID. Only trading days are included, with no entries for weekends or market holidays, leading to regularly spaced but discontinuous sampling.
  • Textual Data: For each trading day, we collect news headlines and article excerpts related to the corresponding company from https://huggingface.co/datasets/Zihan1004/FNSPID. Articles are time-stamped and filtered for relevance. We then use GPT-4.1 Nano to summarize the article into concise 5-sentence narratives, highlighting market developments and company-specific events.
D.5 Resource-Aware Collection (Constraint-Based)
Description
This irregularity arises when data is recorded only while compute resources are active. In this setting, telemetry is missing not due to failure or noise, but because the monitoring system is deliberately inactive during idle periods.
Source
Entity
Each entity corresponds to a distinct physical GPU machine in the Alibaba cluster. We select 3 machines from the dataset, identified by their anonymized worker_name, to represent heterogeneous task schedules and resource usage patterns.
Collection
  • Numerical Time Series: We process eight weeks of Alibaba GPU cluster logs from July–August 2020. For each selected machine, we generate a time series where measurements are sampled only during active periods (i.e., when at least one container is running). At each event timestamp, we log:
    • The number of concurrent instances, tasks, and jobs
    • Eight sensor metrics aggregated across all currently active containers
    The result is an 11-dimensional time series with naturally irregular intervals, reflecting operational constraints and runtime dynamics.
  • Textual Data: For every container start event, we extract the corresponding job_name, task_name, gpu_type_spec, and workload description from the pai_group_tag_table.csv. These fields are treated as an asynchronous textual context. We aggregate all textual entries within a 6-hour window and use GPT-4.1 Nano to summarize them into concise narratives of no more than 5 sentences, capturing key workload patterns and system activity during that period.
D.6 Human Scheduling / Availability (Constraint-Based)
Description
This form of irregularity occurs when data collection is driven by human activity patterns or availability—such as sensors only logging when users are awake or interacting with their devices.
Source
StudentLife study (College Experience Dataset) studentlife .
Entity
Each entity corresponds to an individual participant in the study. We first filter out students with less than 1000 records, then randomly select 20 students from the full cohort to represent a diverse range of behavioral patterns and schedules.
Collection
  • Numerical Time Series: We use smartphone sensor logs—including activity durations on e.g., bike, foot, sleep—recorded at naturally irregular intervals driven by device usage, battery state, and user activity.
  • Textual Data: We extract self-reported survey responses (e.g., anxiety, depression, stress level), each time-stamped by the participant’s device. These serve as asynchronously generated text signals that reflect personal context, mood, or behavior at different points in time.
D.7 Missing Data / Gaps (Artifact-Based)
Description
This type of irregularity results from data that is expected but not recorded, often due to reporting failures, technical issues, or data loss. The result is structured gaps in the time series.
Source
ILINet influenza surveillance (CDC) time_mmd .
Entity
This dataset contains a single entity: the United States as a national-level reporting unit. All observations are aggregated at the country level.
Collection
  • Numerical Time Series: We download weekly counts of influenza-like illness (ILI) cases as reported by the CDC through ILINet. When reporting is incomplete or unavailable for a given week, we preserve these as missing values (NaNs) to reflect true observational gaps.
  • Textual Data: We include weekly CDC public health summaries and provider notes curated from the report and search result collection process introduced by Time-MMD time_mmd . These documents are treated as asynchronous textual observations and provide contextual insight into public health events, seasonal flu trends, and potential reporting disruptions.
D.8 Scheduling Jitter / Delay (Artifact-Based)
Description
This irregularity arises when data points are recorded at uneven intervals due to system delays, congestion, or throttling. Instead of arriving on a regular schedule, observations appear with jitter caused by internal scheduling constraints or resource contention.
Source
CESNET network flow data cesnet .
Entity
Each entity corresponds to a distinct networked device in the infrastructure. We include 11 entities in total, spanning various device types such as IP endpoints, firewalls, routers, servers, and switches.
Collection
  • Numerical Time Series: We ingest flow-level metrics such as byte and packet counts using real-world arrival timestamps. These metrics naturally exhibit jitter due to internal logging delays, variable sampling rates, and system load effects.
  • Textual Data: We collect network device logs that include maintenance messages, traffic control alerts, and system warnings. Each log entry is treated as an independent, asynchronously occurring textual observation, preserving its original timestamp without forcing alignment to numeric data.
D.9 Multi-Source Asynchrony (Artifact-Based)
Description
This type of irregularity arises when multiple data streams—such as sensors—operate with different clocks, sampling rates, or synchronization policies. These asynchronous signals pose challenges for fusion and temporal alignment, especially when modalities are collected independently.
Source
EPA Outdoor Air Quality Sensors epa_air_quality . Numerical time series data was downloaded from: https://aqs.epa.gov/aqsweb/airdata/download_files.html
Textual news articles were retrieved from Common Crawl News, using monthly archives available at: https://data.commoncrawl.org/crawl-data/CC-NEWS/
Entity
Each entity corresponds to a U.S. county monitored by the EPA sensor network. We select 8 counties for inclusion in this dataset:
  • Los Angeles, CA
  • Maricopa (Phoenix), AZ
  • Philadelphia, PA
  • Bexar (San Antonio), TX
  • Dallas, TX
  • Richmond, VA
  • Hillsborough (Tampa), FL
  • Denver, CO
Collection
  • Numerical Time Series: We extract four environmental variables from the EPA’s air monitoring dataset: AQI, Ozone, PM2.5, and Temperature. Each feature is recorded at its native sensor-specific frequency, resulting in asynchronous observations across features and counties.
  • Textual Data: To provide contextual background for environmental conditions, we collect news articles from Common Crawl spanning January to October 2024. Articles are filtered to include only those that mention both the county name and the keyword “weather.” Relevant articles are summarized using GPT-4.1 Nano into concise 5-sentence narratives, highlighting key environmental developments or disruptions affecting the region.
Appendix E Prompt Designed for LLM Summarization
This appendix provides the prompt templates used for GPT-4.1 Nano to generate textual summaries for each dataset in Time-IMM. All prompts are designed to produce concise, factual, and context-aware outputs within a 5-sentence limit.
E.1 GDELT (Event-Based Logging)
PROMPTS = """
You are an expert in {domain}.
This task is part of building the GDELT 2.0 Event dataset.
Below is a news article:
{article}
Write a concise summary of the information related to {domain}, using no more than 5 sentences.
If the article is not relevant to {domain}, reply with exactly: NA
E.2 RepoHealth (Adaptive or Reactive Sampling)
PROMPTS = """
The following are comments from multiple GitHub issue threads in the repository ’{repo_name}’,
including all historical comments accumulated up to and including those made on a specific day.
{comments}
Write a concise summary of the main technical concerns, questions,
or suggestions raised across these discussions. Focus on key issues, proposed solutions,
and unresolved questions. Your summary should be no more than 5 sentences.
E.3 FNSPID (Operational Window Sampling)
PROMPTS = """
You are an expert in financial markets.
This task is part of building the stock price prediction dataset.
Below is a news article:
{article}
Write a concise summary of the information related to financial markets, using no more than 5 sentences.
If the article is not relevant to financial markets, reply with exactly: NA
E.4 ClusterTrace (Resource-Aware Collection)
PROMPTS = """
You are an expert in large-scale distributed computing and cloud infrastructure.
This task is part of building the Alibaba Cluster Trace Program dataset.
Below is a collection of logs and system behavior summaries:
{article}
Write a concise summary of the system behavior and operational insights relevant to distributed computing, using no more than 5 sentences.
If the content is not relevant to the dataset, or the logs/summaries are not provided, reply with exactly: NA
E.5 EPA-Air (Multi-Source Asynchrony)
PROMPTS = """
You are an expert in weather.
This task is part of building the Air Quality Data Collected at Outdoor Monitors Across the US dataset.
Below is a news article:
{article}
Write a concise summary of the information related to weather, using no more than 5 sentences.
If the article is not relevant to weather, reply with exactly: NA
Appendix F Showcase of Raw and Summarized Text
This appendix illustrates how raw textual inputs are preprocessed into concise, informative summaries using LLMs (GPT-4.1 Nano). Below is an example taken from the EPA-Air dataset, where the input news article contains broader event details, but the summarization process emphasizes weather conditions relevant to environmental sensing.
EPA-Air Example
Raw News Excerpt:
Seven migrant smuggling suspects have been arrested and 11 migrants were hospitalized after a law enforcement sting Thursday near San Antonio, authorities said. …
Temperatures in San Antonio were in the high 90s Thursday afternoon and were expected to top 100, according to the National Weather Service.
…The incident came two days after President Joe Biden unveiled plans to enact immediate significant restrictions on migrants seeking asylum …
San Antonio was the site of the nation’s deadliest human smuggling episode in June 2022 …
Summarized Version:
On Thursday in San Antonio, temperatures reached the high 90s and were expected to exceed 100 degrees Fahrenheit. These hot conditions contributed to heat-related injuries among migrants found in a trailer during a law enforcement operation. The extreme heat poses a significant risk for dehydration, heat exhaustion, and other heat-related health issues. Past incidents, such as the deadly smuggling episode in June 2022, also occurred under similar sweltering conditions. The current weather highlights the dangers faced by individuals exposed to high temperatures in the area.
Appendix G Additional Ethical Considerations and Data Access
Data Sources and Privacy.
All datasets in Time-IMM are derived from publicly available sources with appropriate access permissions. For sensitive domains such as healthcare (e.g., MIMIC-IV), we take additional precautions to ensure data privacy and regulatory compliance. MIMIC-IV is a restricted-access resource hosted on https://physionet.org/content/mimiciv/3.1/. To access the raw files, users must be credentialed, complete the CITI “Data or Specimens Only Research” training, and sign the official data use agreement on PhysioNet. In Time-IMM, we do not redistribute raw data from MIMIC-IV. Instead, we provide preprocessing scripts and instructions to enable authorized users to reproduce our setup. We strictly follow access policies and do not feed MIMIC-IV clinical notes into closed-source LLMs (e.g., GPT-4.1 Nano). This ensures alignment with privacy safeguards and ethical standards for clinical data handling.
Responsible Redistribution.
To comply with upstream licenses and ethical research norms, we do not redistribute raw web-crawled text, GitHub comments, or clinical notes where direct sharing is restricted. Instead, we offer derived summaries, structured metadata, and links or instructions for dataset reconstruction from original public sources.
Open Access and Licensing.
Time-IMM is released under a Creative Commons Attribution 4.0 (CC BY 4.0) license, which permits broad use, redistribution, and adaptation with appropriate citation. Scripts and benchmarks are openly available to ensure reproducibility, transparency, and extensibility by the research community.
Hosting and Maintenance Plan.
All datasets and code are hosted on publicly accessible and stable platforms. The dataset is maintained on Kaggle to facilitate broad usability, while the benchmark library is versioned and published through a persistent code repository. We commit to periodically reviewing contributions and issues reported by the community to ensure continued compatibility and correctness.
Accessing.
The Time-IMM dataset can be accessed at: https://www.kaggle.com/datasets/blacksnail789521/time-imm/data
The accompanying benchmark code and processing tools are available at: https://anonymous.4open.science/r/IMMTSF_NeurIPS2025
Ethical Usage Encouragement.
We strongly encourage responsible usage of Time-IMM and IMM-TSF in research contexts that prioritize fairness, transparency, and real-world safety. Applications involving downstream decision-making—especially in domains such as finance or healthcare—should incorporate appropriate validation and interpretability safeguards.
Figure 7: Overview of the canonical pre-alignment process.
Appendix H Adapting Regular Time Series Models via Canonical Pre-Alignment
To adapt regular time series forecasting models to irregular data, we follow the canonical pre-alignment strategy introduced in t-PatchGNN t_patchgnn , as illustrated in Figure 7. This method transforms irregularly sampled time series into a fixed-length, regularly spaced representation that can be consumed by standard forecasting architectures.
Let $X=\{[(t_{i}^{(n)},x_{i}^{(n)})]_{i=1}^{P_{n}}\}_{n=1}^{N}$ denote the set of past irregular time series observations, where $N$ is the number of variables and $P_{n}$ is the number of past observations for variable $n$. Each observation $x_{i}^{(n)}$ occurs at time $t_{i}^{(n)}\in[T_{\text{start}},T_{\text{cut}}]$. We also define a set of future query times $Q=\{[q_{k}^{(n)}]_{k=1}^{F_{n}}\}_{n=1}^{N}$, with each query $q_{k}^{(n)}\in(T_{\text{cut}},T_{\text{end}}]$.
Step 1: Determine Global Temporal Resolution.
We first scan the entire dataset to determine the global maximum number of past input timestamps and future query timestamps across all training windows. This defines the fixed-length resolution $L$ for both input and output time series, covering the combined past and forecast windows.
Step 2: Temporal Grid Construction, Alignment, and Padding.
For each example, we create a unified, chronologically sorted list of timestamps $\{\tilde{t}_{l}\}_{l=1}^{L}$ that covers both the observation period and the forecast horizon. We construct a matrix $\tilde{X}\in\mathbb{R}^{L\times N}$, where each entry $\tilde{x}_{l}^{(n)}$ is the observed value of variable $n$ at time $\tilde{t}_{l}$, or zero if no observation exists. A binary mask matrix $M\in\{0,1\}^{L\times N}$ is used to indicate whether a value was observed (1) or imputed (0). Both $\tilde{X}$ and $M$ are padded as necessary to reach the globally fixed length $L$, ensuring uniform tensor shapes across all examples.
Step 3: Add Input Features: Timestamp and Mask.
To help the model reason about timing and missingness, we expand the feature dimension from $N$ to $12N+$ 1. Each input row includes: (i) the observed values or zeros, (ii) the binary mask, and (iii) the normalized timestamp $\tilde{t}_{l}$. This preserves all temporal and structural information in a dense, aligned format.
Step 4: Add Query Features.
The forecast queries $Q$ are appended to the end of the input sequence using the same $12N+$ 1 feature schema. Each query timestamp $q_{k}^{(n)}$ is represented by a feature row that includes: (i) zero values for all variable entries (since ground truth is unknown at inference time), (ii) a binary mask set to 0 (indicating no observation—not a forecast target), and (iii) the normalized query timestamp. This mask entry should not be confused with the evaluation-time forecast mask used to compute prediction errors. Including query timestamps in the input allows the model to explicitly condition on the temporal locations of future targets.
Summary.
This pre-alignment pipeline transforms irregular, variable-length time series into padded, fixed-length tensors that include time, observation, and mask features. By appending forecast queries and exposing the full temporal grid, regular forecasting models can be repurposed to handle irregular time series in a structured and query-aware manner.
Appendix I Formal Definitions of TTF and MMF Modules
The detailed architectures of the TTF and MMF modules are illustrated in Figures 8 and 9, respectively. We now present their formal definitions below.
I.1 Timestamp-to-Text Fusion (TTF)
Let $S=\{(\tau_{j},s_{j})\}_{j=1}^{L_{S}}$ denote the sequence of past text observations, where $s_{j}$ is a text snippet observed at timestamp $\tau_{j}\in[T_{\text{start}},T_{\text{cut}}]$. Let $v_{j}\in\mathbb{R}^{d}$ be the embedding of $s_{j}$ produced by a fixed language model. Let $Q=\{t_{k}\}_{k=1}^{T_{f}}$ denote the future forecast query times.
Recency-Weighted Averaging (RecAvg).
This variant computes a time-weighted average of all previous embeddings using a Gaussian kernel centered at the forecast time:
$\alpha_{jk}=\exp\left(-\left(\frac{t_{k}-\tau_{j}}{\sigma}\right)^{2}\right),\quad\text{for }\tau_{j}\leq t_{k}$ (6)
$e_{k}=\frac{\sum_{j=1}^{L_{S}}\alpha_{jk}v_{j}}{\sum_{j=1}^{L_{S}}\alpha_{jk}}\in\mathbb{R}^{d}$ (7)
Figure 8: Detailed architecture of TTF modules.
Figure 9: Detailed architecture of MMF modules.
Time2Vec-Augmented Cross-Attention (T2V-XAttn).
Each embedding is augmented with a Time2Vec representation of its timestamp:
$\phi(\tau_{j})=[\omega_{0}\tau_{j}+b_{0},\;\sin(\omega_{1}\tau_{j}+b_{1}),\ldots,\sin(\omega_{d_{\tau}-1}\tau_{j}+b_{d_{\tau}-1})]$ (8)
$\tilde{v}_{j}=\left[\![v_{j};\phi(\tau_{j})]\!\right]\in\mathbb{R}^{d+d_{\tau}}$ (9)
Let $W_{q}\in\mathbb{R}^{d_{q}\times 1}$ be a learnable forecast query vector. Then the aligned representation at time $t_{k}$ is computed by soft attention:
$a_{jk}$ $=\frac{\exp\left((W_{q}^{\top}\tilde{v}_{j})\right)}{\sum_{\tau_{j^{\prime}}\leq t_{k}}\exp\left((W_{q}^{\top}\tilde{v}_{j^{\prime}})\right)}$ (10)
$e_{k}$ $=\sum_{\tau_{j}\leq t_{k}}a_{jk}\cdot\tilde{v}_{j}$ (11)
I.2 Multimodality Fusion (MMF)
Let $Y^{\text{ts}}\in\mathbb{R}^{T_{f}\times N}$ denote the numerical forecast sequence, and let $E=\{e_{k}\}_{k=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times d}$ be the aligned text representations from TTF.
GRU-Gated Residual Addition (GR-Add).
Let $H\in\mathbb{R}^{T_{f}\times h}$ be the hidden state sequence from a GRU applied over the concatenated input:
$z_{k}$ $=\left[\![y_{k}^{\text{ts}};e_{k}]\!\right]\in\mathbb{R}^{N+d}$ (12)
$H$ $=\text{GRU}(\{z_{k}\}_{k=1}^{T_{f}})$ (13)
$\Delta Y$ $=W_{\Delta}H+b_{\Delta}$ (14)
$G$ $=\sigma(W_{g}Z+b_{g})\quad\text{where }Z=[\![Y^{\text{ts}};E]\!]$ (15)
$Y^{\text{fused}}$ $=G\odot Y^{\text{ts}}+(1-G)\odot(Y^{\text{ts}}+\Delta Y)$ (16)
where $W_{\Delta}\in\mathbb{R}^{N\times h}$, $W_{g}\in\mathbb{R}^{N\times(N+d)}$, and $\odot$ denotes element-wise multiplication.
Cross-Attention Addition (XAttn-Add).
Let $Q=Y^{\text{ts}}W_{Q}$, $K=EW_{K}$, and $V=EW_{V}$ with learnable projection matrices $W_{Q},W_{K},W_{V}$. The residual is derived via scaled dot-product attention:
$A$ $=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)\cdot V$ (17)
$\Delta Y$ $=AW_{\text{res}}+b_{\text{res}}$ (18)
$Y^{\text{fused}}$ $=\frac{Y^{\text{ts}}+\kappa\cdot\Delta Y}{1+\kappa}$ (19)
Here, $\kappa\in\mathbb{R}$ is a scalar hyperparameter, and all terms are defined over the entire sequence of query times.
Appendix J Dataset-Specific Forecasting Configuration
Each dataset in Time-IMM is configured with a dataset-specific forecasting setup that reflects its native timestamp distribution and sampling behavior. For each entity, we apply a sliding window strategy with fixed-length context and query segments. The context window corresponds to past observations $X\in[T_{\text{start}},T_{\text{cut}}]$, while the query horizon defines the future timestamps $Q\in(T_{\text{cut}},T_{\text{end}}]$ at which the model is expected to make predictions. Durations are defined using dataset-native time units (e.g., hours, days, or weeks), ensuring semantic consistency across domains.
Table 2: Dataset-specific forecasting parameters: context window (past input) and query horizon (forecast target).
Dataset Context Window Query Horizon
GDELT 14 days 14 days
RepoHealth 31 days 31 days
MIMIC 24 hours 24 hours
FNSPID 31 days 31 days
ClusterTrace 12 hours 12 hours
StudentLife 31 days 31 days
ILINet 4 weeks 4 weeks
CESNET 7 days 7 days
EPA-Air 7 days 7 days
Appendix K Model-Specific and Fusion Module Configuration
K.1 Model-Specific Configuration
Each baseline model is configured using recommended default hyperparameters from its original implementation. The following settings summarize key architecture and training parameters for each model class:
  • Informer informer : 2 encoder layers, 1 decoder layer, attention factor 3.
  • DLinear dlinear : Default configuration without modification.
  • PatchTST patchtst : 1 encoder layer, 1 decoder layer, 2 attention heads.
  • TimesNet timesnet : 2 encoder layers, 1 decoder layer, top-$k=5$, $d_{\text{model}}=16$, $d_{\text{ff}}=32$.
  • TimeMixer timemixer : 2 encoder layers, $d_{\text{model}}=16$, $d_{\text{ff}}=32$, with 3 downsampling layers (window size 2, average pooling).
  • TimeLLM timellm : GPT2 backbone with 6 layers, input/output token lengths 16 and 96, respectively; $d_{\text{model}}=32$, $d_{\text{ff}}=128$.
  • TTM ttm : 3-level attention patching, 3 encoder and 2 decoder layers; $d_{\text{model}}=1024$, decoder dimension 64.
  • CRU cru : Latent state and hidden units set to 32, with square/exponential activations and gravity gates enabled.
  • Latent-ODE latent_ode : Recognition GRU with 32 hidden units, 1 RNN and 1 generator layer, ODE units set to 32.
  • Neural Flow neural_flows : Latent dimension 20, 3-layer hidden network (dimension 32), GRU units 32, coupling flow with 2 layers, time network: TimeLinear.
  • tPatchGNN t_patchgnn : Patch size 24, 1 Transformer layer, 1 GNN layer, 10-dimensional time and node embeddings, $d_{\text{hidden}}=32$.
K.2 Fusion Module Configuration
In multimodal experiments, we integrate textual context into the forecasting pipeline using specialized fusion modules. Each model uses the best-performing fusion configuration selected via validation performance.
Timestamp-to-Text Fusion (TTF).
  • RecAvg: Recency-weighted averaging of past text embeddings using a Gaussian decay function based on timestamp proximity. The maximum weight is assigned when the text timestamp exactly matches the forecast time. We set the standard deviation parameter to $\sigma=1.0$.
  • T2V-XAttn: Each text embedding is augmented with a Time2Vec encoding of its timestamp and fused via a single-layer, single-head attention mechanism.
Multimodality Fusion (MMF).
  • GR-Add: A residual correction is computed from the concatenated numerical forecast and text embedding using a GRU followed by an MLP. The output is blended with the base forecast through a learned gating function.
  • XAttn-Add: Numerical forecasts attend to aligned text representations using a single-layer, single-head attention module. The result is scaled by a fixed convex mixing weight $\kappa=0.5$ and added to the base forecast.
Text Encoder.
We utilize four kinds of pretrained large language models as text encoders—GPT-2, BERT-base, LLaMA 3.1 (8B), and DeepSeek (7B)—each used with its full transformer stack and no truncation of layers. BERT-base supports a maximum context length of 512 tokens, while GPT-2, LLaMA 3.1, and DeepSeek allow longer sequences but are forcibly truncated to 1024 tokens in our setup. To ensure consistent dimensionality, we project all encoder outputs to a shared 768-dimensional space using a learned linear transformation before fusion.
Appendix L More Experimental Results
This appendix contains two sets of additional results that further validate the effectiveness of multimodal modeling in irregular time series forecasting.
  • Section L.1 investigates the role of the text encoder by replacing large language models (LLMs) with a classical alternative—Doc2Vec doc2vec —and comparing the resulting performance.
  • Section L.2 reports per-dataset comparisons between unimodal (time series only) and multimodal (time series + text) model variants across all nine datasets in Time-IMM.
Figure 10: Performance comparison between frozen LLM-based text encoders and Doc2Vec, averaged across all 11 baselines for each dataset.
L.1 Ablation: Replacing LLMs with Doc2Vec
To evaluate the importance of using pretrained language models, we conduct an ablation study replacing frozen LLM-based text encoders with a train-from-scratch alternative—Doc2Vec. Doc2Vec learns fixed-length embeddings for each text input by training a shallow neural model directly on the textual corpus of each dataset, without relying on large-scale pretraining.
Figure 10 summarizes performance across all datasets. We observe that models using frozen LLMs consistently outperform those using Doc2Vec, confirming the benefit of pretrained semantic representations—even in the presence of asynchronous and domain-specific text. These results highlight the importance of strong general-purpose encoders for effective multimodal forecasting under irregular sampling.
L.2 Multimodal vs. Unimodal Performance
To quantify the contribution of textual signals under irregular sampling, we report per-dataset comparisons between unimodal (time series only) and multimodal (time series + text) model variants. Each result reflects the best-performing fusion configuration selected via validation on the corresponding dataset. Performance is measured using mean squared error (MSE), and improvements are reported as relative percentage change from the unimodal baseline.
Multimodal models consistently outperform their unimodal counterparts across all datasets, with especially large improvements observed in settings where text provides rich, temporally relevant context. The following tables summarize detailed results for each dataset:
  • GDELT (Event-Based Logging) — Table 3
  • RepoHealth (Adaptive Sampling) — Table 4
  • MIMIC (Human-Initiated Observation) — Table 5
  • FNSPID (Operational Window Sampling) — Table 6
  • ClusterTrace (Resource-Aware Collection) — Table 7
  • StudentLife (Human Scheduling) — Table 8
  • ILINet (Missing Data) — Table 9
  • CESNET (Scheduling Jitter) — Table 10
  • EPA-Air (Multi-Source Asynchrony) — Table 11
Table 3: Unimodal vs. multimodal forecasting results on the GDELT dataset.
Model Modal MSE MAE
Informer Uni $1.0448$ $0.6935$
Multi $1.0295$ $0.6897$
DLinear Uni $1.0482$ $0.6943$
Multi $1.0487$ $0.6945$
PatchTST Uni $1.0670$ $0.7051$
Multi $1.0511$ $0.6944$
TimesNet Uni $1.0251$ $0.6846$
Multi $1.0250$ $0.6916$
TimeMixer Uni $1.0471$ $0.6842$
Multi $1.0333$ $0.6973$
TimeLLM Uni $1.0411$ $0.6923$
Multi $1.0402$ $0.6894$
TTM Uni $1.0400$ $0.6948$
Multi $1.0238$ $0.6893$
CRU Uni $1.0282$ $0.6968$
Multi $1.0267$ $0.6912$
LatentODE Uni $1.0271$ $0.6886$
Multi $1.0268$ $0.6856$
NeuralFlow Uni $1.0265$ $0.6942$
Multi $1.0260$ $0.6958$
tPatchGNN Uni $1.0267$ $0.6893$
Multi $1.0265$ $0.6958$
Table 4: Unimodal vs. multimodal forecasting results on the RepoHealth dataset.
Model Modal MSE MAE
Informer Uni $0.6391$ $0.4143$
Multi $0.6296$ $0.4037$
DLinear Uni $0.5376$ $0.4081$
Multi $0.5291$ $0.4021$
PatchTST Uni $0.5727$ $0.4185$
Multi $0.5520$ $0.4180$
TimesNet Uni $0.5415$ $0.4158$
Multi $0.5599$ $0.4275$
TimeMixer Uni $0.7267$ $0.4321$
Multi $0.6890$ $0.4230$
TimeLLM Uni $0.5655$ $0.3994$
Multi $0.5601$ $0.3919$
TTM Uni $0.5442$ $0.4088$
Multi $0.5386$ $0.4079$
CRU Uni $1.0362$ $0.8030$
Multi $0.8538$ $0.6596$
LatentODE Uni $0.7555$ $0.6104$
Multi $0.7341$ $0.6285$
NeuralFlow Uni $0.8231$ $0.6736$
Multi $0.7720$ $0.6409$
tPatchGNN Uni $0.6196$ $0.5000$
Multi $0.6090$ $0.4718$
Table 5: Unimodal vs. multimodal forecasting results on the MIMIC dataset.
Model Modal MSE MAE
Informer Uni $0.8519$ $0.5588$
Multi $0.7988$ $0.6083$
DLinear Uni $0.8770$ $0.6152$
Multi $0.8515$ $0.6456$
PatchTST Uni $0.9394$ $0.6794$
Multi $0.9281$ $0.6799$
TimesNet Uni $1.0374$ $0.7348$
Multi $0.9335$ $0.6821$
TimeMixer Uni $0.8856$ $0.5686$
Multi $0.8179$ $0.5821$
TimeLLM Uni $0.7901$ $0.5457$
Multi $0.7535$ $0.5620$
TTM Uni $0.8620$ $0.6283$
Multi $0.9105$ $0.6707$
CRU Uni $1.0122$ $0.7235$
Multi $0.9847$ $0.7087$
LatentODE Uni $0.9325$ $0.6868$
Multi $0.9423$ $0.6785$
NeuralFlow Uni $1.0086$ $0.7219$
Multi $0.9884$ $0.7079$
tPatchGNN Uni $0.7250$ $0.5479$
Multi $0.7656$ $0.5794$
Table 6: Unimodal vs. multimodal forecasting results on the FNSPID dataset.
Model Modal MSE MAE
Informer Uni $0.1277$ $0.2153$
Multi $0.1282$ $0.2149$
DLinear Uni $0.1239$ $0.2121$
Multi $0.1248$ $0.2145$
PatchTST Uni $0.1301$ $0.2240$
Multi $0.1277$ $0.2212$
TimesNet Uni $0.1253$ $0.2144$
Multi $0.1239$ $0.2128$
TimeMixer Uni $0.1161$ $0.1965$
Multi $0.1157$ $0.1964$
TimeLLM Uni $0.1261$ $0.2138$
Multi $0.1240$ $0.2124$
TTM Uni $0.1355$ $0.2373$
Multi $0.1335$ $0.2312$
CRU Uni $0.1484$ $0.2545$
Multi $0.1476$ $0.2471$
LatentODE Uni $0.1203$ $0.2134$
Multi $0.1157$ $0.2051$
NeuralFlow Uni $0.1510$ $0.2572$
Multi $0.1420$ $0.2470$
tPatchGNN Uni $0.1247$ $0.2195$
Multi $0.1212$ $0.2132$
Table 7: Unimodal vs. multimodal forecasting results on the ClusterTrace dataset.
Model Modal MSE MAE
Informer Uni $1.0793$ $0.8039$
Multi $0.7473$ $0.6527$
DLinear Uni $0.8418$ $0.6765$
Multi $0.6440$ $0.6504$
PatchTST Uni $2.0366$ $0.9896$
Multi $1.0370$ $0.8389$
TimesNet Uni $0.9739$ $0.7187$
Multi $0.6706$ $0.6579$
TimeMixer Uni $1.1262$ $0.7861$
Multi $0.7377$ $0.6946$
TimeLLM Uni $0.9596$ $0.7325$
Multi $0.7250$ $0.6973$
TTM Uni $1.0030$ $0.7409$
Multi $0.7166$ $0.6813$
CRU Uni $1.2111$ $0.8287$
Multi $0.5753$ $0.5988$
LatentODE Uni $1.1613$ $0.7818$
Multi $0.5336$ $0.5780$
NeuralFlow Uni $1.4458$ $0.8910$
Multi $0.6248$ $0.6384$
tPatchGNN Uni $0.9671$ $0.7119$
Multi $0.6105$ $0.6104$
Table 8: Unimodal vs. multimodal forecasting results on the StudentLife dataset.
Model Modal MSE MAE
Informer Uni $0.9432$ $0.6674$
Multi $0.8910$ $0.6790$
DLinear Uni $0.8850$ $0.6760$
Multi $0.8733$ $0.6629$
PatchTST Uni $0.8964$ $0.6793$
Multi $0.8764$ $0.6814$
TimesNet Uni $0.9069$ $0.6827$
Multi $0.8892$ $0.6889$
TimeMixer Uni $0.9437$ $0.6720$
Multi $0.8867$ $0.6768$
TimeLLM Uni $0.9355$ $0.6797$
Multi $0.8870$ $0.6781$
TTM Uni $0.9024$ $0.6830$
Multi $0.8888$ $0.6838$
CRU Uni $0.9101$ $0.6876$
Multi $0.9101$ $0.6890$
LatentODE Uni $0.9028$ $0.6965$
Multi $0.9040$ $0.6907$
NeuralFlow Uni $0.9237$ $0.6995$
Multi $0.9150$ $0.7024$
tPatchGNN Uni $0.8661$ $0.6672$
Multi $0.8633$ $0.6705$
Table 9: Unimodal vs. multimodal forecasting results on the ILINet dataset.
Model Modal MSE MAE
Informer Uni $1.2714$ $0.8118$
Multi $1.2338$ $0.7832$
DLinear Uni $1.0402$ $0.6811$
Multi $0.9795$ $0.6374$
PatchTST Uni $1.1606$ $0.7532$
Multi $1.1727$ $0.7373$
TimesNet Uni $1.0518$ $0.6670$
Multi $1.0510$ $0.6436$
TimeMixer Uni $0.9825$ $0.6456$
Multi $1.0530$ $0.6845$
TimeLLM Uni $1.1243$ $0.7255$
Multi $1.0333$ $0.6901$
TTM Uni $1.1010$ $0.7003$
Multi $1.0383$ $0.6898$
CRU Uni $0.9718$ $0.6652$
Multi $1.0423$ $0.6900$
LatentODE Uni $1.1512$ $0.7531$
Multi $1.1098$ $0.7276$
NeuralFlow Uni $1.1946$ $0.7688$
Multi $1.1501$ $0.7460$
tPatchGNN Uni $1.6163$ $0.9255$
Multi $1.4877$ $0.8689$
Table 10: Unimodal vs. multimodal forecasting results on the CESNET dataset.
Model Modal MSE MAE
Informer Uni $0.9591$ $0.7559$
Multi $0.8754$ $0.7233$
DLinear Uni $0.9695$ $0.7517$
Multi $0.8763$ $0.7221$
PatchTST Uni $1.3177$ $0.8797$
Multi $1.1242$ $0.8192$
TimesNet Uni $0.9581$ $0.7499$
Multi $0.8686$ $0.7201$
TimeMixer Uni $0.9500$ $0.7481$
Multi $0.8588$ $0.7160$
TimeLLM Uni $0.9891$ $0.7676$
Multi $0.9125$ $0.7364$
TTM Uni $0.9869$ $0.7653$
Multi $0.9205$ $0.7422$
CRU Uni $0.9462$ $0.7559$
Multi $0.8726$ $0.7257$
LatentODE Uni $0.9655$ $0.7642$
Multi $0.8977$ $0.7400$
NeuralFlow Uni $1.0154$ $0.7915$
Multi $0.9495$ $0.7685$
tPatchGNN Uni $0.9796$ $0.7713$
Multi $0.9122$ $0.7483$
Table 11: Unimodal vs. multimodal forecasting results on the EPA-Air dataset.
Model Modal MSE MAE
Informer Uni $0.6301$ $0.5983$
Multi $0.5812$ $0.5728$
DLinear Uni $0.5361$ $0.5279$
Multi $0.5223$ $0.5223$
PatchTST Uni $0.6196$ $0.5947$
Multi $0.6204$ $0.5992$
TimesNet Uni $0.5599$ $0.5524$
Multi $0.5892$ $0.5650$
TimeMixer Uni $0.6086$ $0.5770$
Multi $0.5641$ $0.5663$
TimeLLM Uni $0.5835$ $0.5643$
Multi $0.5334$ $0.5303$
TTM Uni $0.6002$ $0.5653$
Multi $0.6218$ $0.5761$
CRU Uni $0.7026$ $0.6306$
Multi $0.7982$ $0.6739$
LatentODE Uni $0.8025$ $0.6665$
Multi $0.7556$ $0.6523$
NeuralFlow Uni $0.7821$ $0.6488$
Multi $0.8202$ $0.6790$
tPatchGNN Uni $0.6258$ $0.6022$
Multi $0.5840$ $0.5793$